Which number exceeds the difference between 300 and 100 by 500?
Answer by ikleyn(52921)

.
Which number exceeds the difference between 300 and 100 by 500?
~~~~~~~~~~~~~~~~~~~~~~~~

    This number is  (300-100) + 500 = 200 + 500 = 700.    ANSWER

Solved.